# arXiv:2603.01932v1[cs.CV] 02 Mar 2026

### v1

@@ 3.1.2 Flight Mission Geometry and Coverage
3.1.3 Image Preprocessing
Raw multispectral frames are converted to single precision arrays, filtered by capture quality flags, and denoised with a $3\times 3$ median kernel applied independently to the R, G, B, NIR and RedEdge bands. Frames that contain more than 0.5% saturated pixels in any band are discarded, and lens vignetting and distortion are corrected using the vendor calibration so that all subsequent steps operate on rectified images. Cross band alignment then registers every band to the green reference. SIFT keypoints are extracted per band, descriptor matches are filtered with a Lowe ratio test and RANSAC, and a projective homography is estimated for each band, which is adequate for the near planar scenes at 120 m altitude. The transform is refined by maximizing the enhanced correlation coefficient on a grid of small windows and registration quality is reported as the median reprojection error, with frames above a 0.5 pixel threshold flagged for reprocessing. Figure 3 illustrates typical inlier correspondences after matching and RANSAC.
Panel based radiometric calibration converts the registered digital numbers to per band reflectance that is consistent across sorties. For each band a scale factor is derived from pre and post flight panel captures and the upward irradiance sensor, and per frame reflectance is adjusted for exposure and irradiance variation, clipped to a fixed range, and stored as float32. Orthorectification then uses RTK geotags and structure from motion in Agisoft Metashape to estimate camera poses and a digital elevation model. All bands of each frame are treated as a rigid group and are projected onto the elevation model into a common ground coordinate system. Seamline optimization and blending within the overlap generate multispectral and RGB orthomosaics without exposure compensation for the analytical bands so that radiometric consistency is preserved. The final per band mosaics are exported as tiled GeoTIFFs for index computation and patch extraction in the dataset pipeline.
Figure 5: Overview of the proposed pipeline. Bottom: SRAB radiance stream with residual attention and a U-Net decoder. Top: vegetation-index stream with windowed self-attention, Mamba blocks, Slot Attention with mean-slot broadcast, and a scale-aligned decoder. Right: fusion head that concatenates the two 64-channel feature maps at the native resolution to generate class logits.
3.1.4 Annotation and Dataset Finalization
Pixels in the multispectral orthomosaics are labelled at native resolution as barley crop, ryegrass weed or other, where other covers soil, wheel tracks and infrastructure. Orthorectified and radiometrically normalized mosaics are tiled, and a vegetation mask derived from NDVI restricts annotation to canopy pixels while nonvegetation and invalid pixels are assigned other or an ignore value. Within this mask annotators use a GIS editor with orthomosaic, NDVI and RGB composites to draw crop and weed polygons, and very small polygons are discarded. The polygons are rasterised by assigning each pixel the class of the polygon that contains its centre, overlaps are resolved in favour of weed polygons so that mixed canopies follow the weed label, and remaining vegetation pixels outside all polygons are kept as other. Each tile is annotated by one operator and checked by a second, and a stratified subset is reannotated to maintain consistent class usage across fields and years. The final labels are stored as single channel GeoTIFFs with other equal to zero, crop equal to one, weed equal to two and ignore equal to two hundred and fifty five, and during training this map is converted to one hot targets $y_{b,c,h,w}\in\{0,1\}$ with the ignore value forming a binary mask that removes those pixels from the supervised loss.
@@ 3.2 Methods
We propose VISA, a two-stream segmentation architecture that leverages complementary cues from vegetation indices and multispectral imagery. The model aims to jointly capture fine spatial detail and field-scale agronomic patterns that are difficult to resolve from a single feature stream. As shown in Figure 5, the network contains two parallel branches, a Vegetation-Index Branch that operates on normalized vegetation indices and a Spectral Residual Attention Branch (SRAB) that processes the multispectral radiance cube. Both branches output 64-channel feature maps at native $256\times 256$ resolution, which are fused by a shallow head to produce per-pixel logits for crop, weed, and background classes. This design separates low-level radiometric reasoning from high-level vegetation context modeling, improving robustness under illumination changes and canopy mixing. Let $B$ denote batch size. Each training sample is a $256\times 256$ patch cropped from a georeferenced orthomosaic. The radiance input is a five-band reflectance tensor $X^{\mathrm{raw}}\in\mathbb{R}^{B\times 5\times H\times W}$ with H=W=256H{=}W{=}256, ordered as $(B,G,R,\mathrm{RE},\mathrm{NIR})$ after band co-registration. The index input is a five-channel tensor $X^{\mathrm{idx}}\in\mathbb{R}^{B\times 5\times H\times W}$ formed by vegetation index maps computed from the same reflectance bands. Unless otherwise stated, convolutions use zero padding and preserve spatial size. Both branches output native-resolution features in $\mathbb{R}^{B\times 64\times 256\times 256}$ and are fused to predict C=3C{=}3 semantic classes $\{\text{other},\text{crop},\text{weed}\}$, while the ignore label is excluded from supervision by a binary mask. The configuration used throughout the paper fixes the embedding width at d=64d{=}64, the window size at s=8s{=}8 for windowed self-attention, the number of Slot Attention groups at K=6K{=}6, and the number of Slot Attention refinement iterations at T=3T{=}3.
3.2.1 Vegetation-Index Modeling Branch
@@
To complement raw-band learning with normalized spectral reasoning, we introduce a vegetation-index modeling branch that operates on a compact stack of five indices $\{\mathrm{NDVI},\mathrm{GNDVI},\\ \mathrm{EVI},\mathrm{SAVI},\mathrm{MSAVI}\}$ derived from calibrated reflectance. These indices emphasize vegetation vigor and suppress illumination variability, allowing the model to capture agronomic patterns that arise from cross-band interactions rather than single-band responses. Indices are computed from calibrated reflectance with a small stabilizer $\epsilon{=}10^{-6}$. For example, $\mathrm{NDVI}=(\mathrm{NIR}-R)/(\mathrm{NIR}+R+\epsilon)$, $\mathrm{GNDVI}=(\mathrm{NIR}-G)/(\mathrm{NIR}+G+\epsilon)$, and $\mathrm{SAVI}=1.5(\mathrm{NIR}-R)/(\mathrm{NIR}+R+0.5+\epsilon)$. We apply a fixed per-index standardization using statistics computed on the training split only. For the $i$th index channel, $\tilde{X}^{\mathrm{idx}}_{i}=(X^{\mathrm{idx}}_{i}-\mu_{i})/(\sigma_{i}+\epsilon)$, and the same $(\mu_{i},\sigma_{i})$ are reused for validation and test data to avoid leakage.
@@
Each index map is linearly standardized and projected to an embedding width d=64d{=}64 via a $1{\times}1$ convolution and LayerNorm:
@@
Here $W^{\mathrm{proj}}\in\mathbb{R}^{d\times 5\times 1\times 1}$ and $b^{\mathrm{proj}}\in\mathbb{R}^{d}$ implement a $1\times 1$ projection from five indices to d=64d{=}64 channels. $\text{LN}(\cdot)$ denotes LayerNorm applied over the channel dimension at each spatial location, which stabilizes training under index magnitude variations across seasons and fields.
We reshape $Z$ into non-overlapping $8{\times}8$ windows and apply two lightweight Transformer encoder layers with windowed self-attention. This design preserves local spatial coherence (e.g., crop rows and patch textures) while avoiding the quadratic complexity of global attention. A decomposed relative positional bias further enhances geometric sensitivity within each window.
@@
where $\mathrm{WSA}$ denotes the multi-head windowed attention above. $\mathrm{FFN}$ is a position-wise two-layer perceptron with GELU nonlinearity applied independently to each token. In our configuration, s=8s{=}8 and h=8h{=}8, which keeps attention local to $64$ tokens per window while preserving row-level texture cues.
@@
We apply the state-space blocks on a 1D token sequence obtained by raster scanning the $H\times W$ grid, giving L=HWL{=}HW tokens. At step $t$, $u_{t}\in\mathbb{R}^{d}$ is the token embedding and $x_{t}\in\mathbb{R}^{d}$ is the state vector. $\bar{A}\in\mathbb{R}^{d}$ and $\bar{B}\in\mathbb{R}^{d}$ are learned per-channel coefficients after discretization, and $\odot$ denotes element-wise multiplication. The gate $\sigma(W_{g}u_{t})$ uses $W_{g}\in\mathbb{R}^{d\times d}$ and modulates input injection to suppress ambiguous mixed pixels. This recurrence updates each token in linear time in $L$ and does not incur the quadratic cost of global attention, which is important for field-scale context modeling on dense orthomosaics.
Finally, we apply Slot Attention (K=6K{=}6 slots, T=3T{=}3 iterations) to group tokens into structured semantic regions (e.g., crop canopy, weed clusters, soil). The aggregated descriptor is broadcasted back to the spatial grid, forming a feature map $F^{\mathrm{idx}}\in\mathbb{R}^{B\times 64\times 256\times 256}$ that carries globally consistent vegetation cues. Let $U^{+}\in\mathbb{R}^{L\times d}$ denote the token matrix after windowed attention and state-space filtering. Slot Attention maintains $K$ slot vectors $S=\{s_{k}\}_{k=1}^{K}$ with $s_{k}\in\mathbb{R}^{d}$. At each refinement iteration, we compute attention weights between tokens and slots by
@@
This feature map is decoded by shallow convolutional upsampling layers and supervised by an auxiliary segmentation loss weighted by $0.3$, encouraging early-stage discriminative learning in the index domain.
@@ 3.2.2 Spectral Residual Attention Branch
In parallel, the spectral residual attention branch learns directly from calibrated multispectral reflectance $\{B,G,R,\mathrm{RE},\mathrm{NIR}\}$ to retain subtle spectral contrasts that ratio-based indices may suppress. The branch adopts a three-stage residual-attention encoder followed by a symmetric decoder.
Each stage stacks two residual–attention units consisting of convolutional bodies, squeeze–excitation channel weighting, and CBAM-style spatial gating. Formally, a unit transforms feature $U$ as
@@
The squeeze-excitation operator produces channel weights from global statistics. Let $g=\frac{1}{HW}\sum_{h,w}U_{:,h,w}\in\mathbb{R}^{C}$ be global average pooled features for $C$ channels. SE computes $a=\sigma\!\left(W_{2}^{\mathrm{se}}\mathrm{GELU}(W_{1}^{\mathrm{se}}g)\right)$ and rescales channels by $U^{\mathrm{se}}_{c,h,w}=a_{c}\,U_{c,h,w}$, where $W_{1}^{\mathrm{se}}$ and $W_{2}^{\mathrm{se}}$ are learned linear layers and $\sigma(\cdot)$ is the sigmoid.
@@
Encoded features with widths $\{64,128,256\}$ are hierarchically aggregated by transposed convolutions and refinement blocks, restoring native $256{\times}256$ resolution. The resulting feature map $F^{\mathrm{raw}}\!\in\!\mathbb{R}^{B\times 64\times 256\times 256}$ encodes high-frequency appearance cues complementary to the vegetation-index pathway.
The encoder uses three resolution levels with channel widths $\{64,128,256\}$. Spatial resolution is reduced by a factor of two between levels, and the decoder restores $256\times 256$ resolution with learnable upsampling. Skip connections from the three encoder levels are concatenated into the corresponding decoder stages to preserve row boundaries and small weed fragments that are lost under repeated downsampling. The last decoder stage outputs $F^{\mathrm{raw}}$ at width $64$ to match the index branch for fusion.
@@ 3.2.3 Feature Fusion and Prediction Head
Both feature maps are aligned on the same $256\times 256$ grid and share the same channel width, which allows fusion without interpolation. The classifier predicts C=3C{=}3 logits per pixel and the ignore label is implemented only in the supervision mask rather than as a prediction class.
Let $Y^{\mathrm{raw}},Y^{\mathrm{idx}}\in\mathbb{R}^{B\times 64\times 256\times 256}$ be the branch outputs at native resolution. They are concatenated along the channel axis to form $F^{\mathrm{cat}}=\mathrm{Concat}(Y^{\mathrm{raw}},Y^{\mathrm{idx}})\in\mathbb{R}^{B\times 128\times 256\times 256}$. The fusion block first performs channel mixing and local aggregation by a $3{\times}3$ convolution with zero padding, followed by batch normalization and a rectified linear activation. Denote the fusion kernel $W^{\mathrm{fus}}\in\mathbb{R}^{64\times 128\times 3\times 3}$ and bias $b^{\mathrm{fus}}\in\mathbb{R}^{64}$. The pre-activation fused map $F\in\mathbb{R}^{B\times 64\times 256\times 256}$ is defined component wise as
@@
Training minimizes a class-balanced cross-entropy on the fused logits with an ignore mask for unlabeled pixels. Let $\Omega$ be the set of labeled pixels over a mini-batch and $y_{b,c,h,w}\in\{0,1\}$ the one-hot targets. Class weights $w_{c}$ follow median-frequency balancing with pixel frequencies $f_{c}$ measured on the training set, that is $w_{c}=\mathrm{median}(f_{1},\ldots,f_{C})/f_{c}$. The primary loss is
@@
To improve robustness to class imbalance and to sharpen thin weed boundaries, we add a soft Dice term and an edge-aware term computed on the labeled pixels. Let $P_{b,c,h,w}$ be the fused posterior in (11) and $y_{b,c,h,w}$ the one-hot target. The Dice loss is

### v2

@@ 3.2 Methods
We propose VISA, a two-stream segmentation architecture that leverages complementary cues from vegetation indices and multispectral imagery. The model aims to jointly capture fine spatial detail and field-scale agronomic patterns that are difficult to resolve from a single feature stream. As shown in Figure 5, the network contains two parallel branches, a Vegetation Index Branch that operates on normalized vegetation indices and a Spectral Residual Attention Branch (SRAB) that processes the multispectral radiance cube. Both branches output 64-channel feature maps at native $256\times 256$ resolution, which are fused by a shallow head to produce per-pixel logits for crop, weed, and background classes. This design separates low-level radiometric reasoning from high-level vegetation context modelling, improving robustness under illumination changes and canopy mixing. Let $B$ denote batch size. Each training sample is a $256\times 256$ patch cropped from a georeferenced orthomosaic. The radiance input is a five-band reflectance tensor $X^{\mathrm{raw}}\in\mathbb{R}^{B\times 5\times H\times W}$ with H=W=256H~{=}~W~{=}~256, ordered as $(B,G,R,\mathrm{RE},\mathrm{NIR})$ after band co-registration. The index input is a five-channel tensor $X^{\mathrm{idx}}\in\mathbb{R}^{B\times 5\times H\times W}$ formed by vegetation-index maps computed from the same reflectance bands. Unless otherwise stated, convolutions use zero padding and preserve spatial size. Both branches output native-resolution features in $\mathbb{R}^{B\times 64\times 256\times 256}$ and are fused to predict C=3C~{=}~3 semantic classes $\{\text{other},\text{crop},\text{weed}\}$, while the ignore label is excluded from supervision by a binary mask. The configuration used throughout the paper fixes the embedding width at d=64d~{=}~64, the window size at s=8s~{=}~8 for windowed self-attention, the number of Slot Attention groups at K=6K~{=}~6, and the number of Slot Attention refinement iterations at T=3T~{=}~3.
3.2.1 Vegetation-Index Modelling Branch
@@
To complement raw-band learning with normalized spectral reasoning, we introduce a vegetation-index modelling branch that operates on a compact stack of five indices $\{\mathrm{NDVI},\mathrm{GNDVI},\mathrm{EVI},\mathrm{SAVI},\mathrm{MSAVI}\}$ derived from calibrated reflectance. These five indices were selected to provide complementary responses within the five-band sensing capability of the DJI Phantom 4 Multispectral. NDVI serves as a standard vegetation-vigor baseline. GNDVI increases sensitivity to chlorophyll variation through the green band. EVI incorporates the blue band and is less sensitive to canopy background and saturation at higher vegetation density. SAVI and MSAVI explicitly reduce soil-background effects, which is relevant in BAWSeg because the four-year barley scenes include both partially exposed soil and incomplete canopy closure. Therefore, we selected indices that jointly capture canopy vigor, chlorophyll-sensitive variation, reduced saturation, and soil-adjusted vegetation response, rather than relying on a single family of closely related ratios. Indices are computed from calibrated reflectance with a small stabilizer $\epsilon~{=}~10^{-6}$. For example, $\mathrm{NDVI}=(\mathrm{NIR}-R)/(\mathrm{NIR}+R+\epsilon)$, $\mathrm{GNDVI}=(\mathrm{NIR}-G)/(\mathrm{NIR}+G+\epsilon)$, and $\mathrm{SAVI}=1.5(\mathrm{NIR}-R)/(\mathrm{NIR}+R+0.5+\epsilon)$. We apply a fixed per-index standardization using statistics computed on the training split only. For the $i$th index channel, $\tilde{X}^{\mathrm{idx}}_{i}=(X^{\mathrm{idx}}_{i}-\mu_{i})/(\sigma_{i}+\epsilon)$, and the same $(\mu_{i},\sigma_{i})$ are reused for validation and test data to avoid leakage.
Figure 5: Overview of the proposed pipeline. Bottom: SRAB radiance stream with residual attention and a U-Net decoder. Top: vegetation-index stream with windowed self-attention, Mamba blocks, Slot Attention with mean-slot broadcast, and a scale-aligned refinement module. Here, the scale-aligned refinement module denotes the lightweight feature-refinement stage that aligns the index-stream representation with the radiance branch on the native $256\times 256$ grid before fusion. Right: fusion head that concatenates the two 64-channel feature maps at the native resolution to generate class logits.
@@
Each index map is linearly standardized and projected to an embedding width d=64d~{=}~64 via a $1~{\times}~1$ convolution and LayerNorm:
@@
Here, $W^{\mathrm{proj}}\in\mathbb{R}^{d\times 5\times 1\times 1}$ and $b^{\mathrm{proj}}\in\mathbb{R}^{d}$ implement a $1\times 1$ projection from five indices to d=64d~{=}~64 channels. $\text{LN}(\cdot)$ denotes LayerNorm applied over the channel dimension at each spatial location, which stabilizes training under index magnitude variations across seasons and fields.
We reshape $Z$ into non-overlapping $8~{\times}~8$ windows and apply two lightweight Transformer encoder layers with windowed self-attention. This design preserves local spatial coherence (e.g., crop rows and patch textures) while avoiding the quadratic complexity of global attention. A decomposed relative positional bias further enhances geometric sensitivity within each window.
@@
where $\mathrm{WSA}$ denotes the multi-head windowed attention above. $\mathrm{FFN}$ is a position-wise two-layer perceptron with GELU nonlinearity applied independently to each token. In our configuration, s=8s~{=}~8 and h=8h~{=}~8, which keeps attention local to $64$ tokens per window while preserving row-level texture cues.
@@
We apply the state-space blocks on a 1D token sequence obtained by a row-major raster scan of the $H\times W$ grid, giving L=HWL~{=}~HW tokens. The current implementation uses a single scan order only and does not adopt bidirectional, reverse, or cross-scanning variants. At step $t$, $u_{t}\in\mathbb{R}^{d}$ is the token embedding and $x_{t}\in\mathbb{R}^{d}$ is the state vector. $\bar{A}\in\mathbb{R}^{d}$ and $\bar{B}\in\mathbb{R}^{d}$ are learned per-channel coefficients after discretization, and $\odot$ denotes element-wise multiplication. The gate $\sigma(W_{g}u_{t})$ uses $W_{g}\in\mathbb{R}^{d\times d}$ and modulates input injection to suppress ambiguous mixed pixels. In this design, directional bias is partially mitigated because the sequential module operates after windowed self-attention with relative positional bias, while spatial augmentation by rotations and flips is applied during training. This recurrence updates each token in linear time in $L$ and does not incur the quadratic cost of global attention, which is important for field-scale context modelling on dense orthomosaics.
Finally, we apply Slot Attention (K=6K~{=}~6 slots, T=3T~{=}~3 iterations) to group tokens into structured semantic regions (e.g., crop canopy, weed clusters, soil). The aggregated descriptor is broadcasted back to the spatial grid, forming a feature map $F^{\mathrm{idx}}\in\mathbb{R}^{B\times 64\times 256\times 256}$ that carries globally consistent vegetation cues. Let $U^{+}\in\mathbb{R}^{L\times d}$ denote the token matrix after windowed attention and state-space filtering. Slot Attention maintains $K$ slot vectors $S=\{s_{k}\}_{k=1}^{K}$ with $s_{k}\in\mathbb{R}^{d}$. At each refinement iteration, we compute attention weights between tokens and slots by
@@
The resulting feature map is then passed to a scale-aligned refinement module, which consists of lightweight convolutional refinement layers applied on the native $256\times 256$ grid. Its role is to refine the index-stream representation before fusion and to align it with the radiance branch while preserving the contextual information aggregated by the Transformer, Mamba, and Slot Attention modules. The refined index feature map is additionally supervised by an auxiliary segmentation loss weighted by $0.3$, which encourages discriminative learning in the index domain.
@@ 3.2.2 Spectral Residual Attention Branch
In parallel, the spectral residual attention branch learns directly from calibrated multispectral reflectance $\{B,G,R,\mathrm{RE},\mathrm{NIR}\}$ rather than from normalized vegetation-index ratios. This choice preserves local band-specific contrast at native resolution, including weak spectral discontinuities along crop–weed boundaries, fine row texture, and fragmented responses from small weed patches that may be attenuated after ratio-based compression. The branch adopts a three-stage residual-attention encoder followed by a symmetric decoder. Local residual convolutions capture short-range spatial detail, while the attention components reweight informative channels and spatial locations under task supervision.
Each stage stacks two residual-attention units consisting of convolutional bodies, squeeze–excitation channel weighting, and CBAM-style spatial gating. Formally, a unit transforms feature $U$ as
@@
The squeeze–excitation operator produces channel weights from global statistics. Let $g=\frac{1}{HW}\sum_{h,w}U_{:,h,w}\in\mathbb{R}^{C}$ be global average-pooled features for $C$ channels. SE computes $a=\sigma\!\left(W_{2}^{\mathrm{se}}\mathrm{GELU}(W_{1}^{\mathrm{se}}g)\right)$ and rescales channels by $U^{\mathrm{se}}_{c,h,w}=a_{c}\,U_{c,h,w}$, where $W_{1}^{\mathrm{se}}$ and $W_{2}^{\mathrm{se}}$ are learned linear layers and $\sigma(\cdot)$ is the sigmoid.
@@
Encoded features with widths $\{64,128,256\}$ are hierarchically aggregated by transposed convolutions and refinement blocks, restoring the native $256~{\times}~256$ resolution. The resulting feature map $F^{\mathrm{raw}}\!\in\!\mathbb{R}^{B\times 64\times 256\times 256}$ encodes high-frequency appearance cues complementary to the vegetation-index pathway. Skip connections from the encoder to the decoder help preserve boundary localization and reduce the loss of thin or spatially sparse weed structures during downsampling.
The encoder uses three resolution levels with channel widths $\{64,128,256\}$. Spatial resolution is reduced by a factor of two between levels, and the decoder restores the $256\times 256$ resolution with learnable upsampling. Skip connections from the three encoder levels are concatenated into the corresponding decoder stages to preserve row boundaries and small weed fragments that are lost under repeated downsampling. The last decoder stage outputs $F^{\mathrm{raw}}$ at width $64$ to match the index branch for fusion.
@@ 3.2.3 Feature Fusion and Prediction Head
Both feature maps are aligned on the same $256\times 256$ grid and share the same channel width, which allows fusion without interpolation. The classifier predicts C=3C~{=}~3 logits per pixel and the ignore label is implemented only in the supervision mask rather than as a prediction class.
Let $Y^{\mathrm{raw}},Y^{\mathrm{idx}}\in\mathbb{R}^{B\times 64\times 256\times 256}$ be the branch outputs at native resolution. They are concatenated along the channel axis to form $F^{\mathrm{cat}}=\mathrm{Concat}(Y^{\mathrm{raw}},Y^{\mathrm{idx}})\in\mathbb{R}^{B\times 128\times 256\times 256}$. The fusion block first performs channel mixing and local aggregation by a $3~{\times}~3$ convolution with zero padding, followed by batch normalization and a rectified linear activation. Denote the fusion kernel $W^{\mathrm{fus}}\in\mathbb{R}^{64\times 128\times 3\times 3}$ and bias $b^{\mathrm{fus}}\in\mathbb{R}^{64}$. The pre-activation fused map $F\in\mathbb{R}^{B\times 64\times 256\times 256}$ is defined component-wise as
@@
Training minimizes a class-balanced cross-entropy on the fused logits with an ignore mask for unlabeled pixels. Let $\Omega$ be the set of labeled pixels over a mini-batch and $y_{b,c,h,w}\in\{0,1\}$ be the one-hot targets. Class weights $w_{c}$ follow median-frequency balancing with pixel frequencies $f_{c}$ measured on the training set, that is, $w_{c}=\mathrm{median}(f_{1},\ldots,f_{C})/f_{c}$. The primary loss is
@@
To improve robustness to class imbalance and to sharpen thin weed boundaries, we add a soft Dice term and an edge-aware term computed on the labeled pixels. Let $P_{b,c,h,w}$ be the fused posterior in (11) and $y_{b,c,h,w}$ be the one-hot target. The Dice loss is